OpenStudy (anonymous):

what is 5 divided by 4.35

OpenStudy (nechirwan):

what do u mean ?

OpenStudy (anonymous):

5/4.35=1.149 or 1.15 or 4.35/5=0.87

OpenStudy (cuanchi):

5/4.35=1.1494 if you are talking about significant figures the answer will be "1" because 5 has only one significant figure. But if 5 is an exact number, then the answer will be "1.15" because the 4.35 has 3 significant figures and the digit after 4 is >than 4.